OpenStudy (anonymous):

express the limit as a derivative and evaluate. lim cos(theta)-0.5 over theta-pi/3 x approaches pi/3

OpenStudy (anonymous):

\[\lim x >>\pi/3 \frac{ \cos (\theta)-0.5 }{ \theta-\pi/3 }\]